OpenStudy (anonymous):

A drawer holds 8 red socks, 13 blue socks, and 15 green socks. What is the probability that a red sock will be drawn first and then a green sock (without replacement)?

OpenStudy (kropot72):

If we do this in steps, the first step is to find the probability that a red sock will be drawn first. This is given by: \[\large \frac{number\ of\ red\ socks}{total\ number\ of\ all\ colored\ socks}=you\ can\ calculate\] When you have found the probability of a red sock on the first draw, we can do the final step.

OpenStudy (kropot72):

@vbryand Are you there?

OpenStudy (anonymous):

8/36 yes?

OpenStudy (kropot72):

Good work! You are correct. The probability of drawing a green sock on the second draw is given by: \[\large \frac{number\ of\ green\ socks}{total\ remaining\ number\ of\ socks}\] When you have found the probability of a green sock on the second draw, the probability that a red sock will be drawn first and then a green sock is given by: \[P(red\ on\ first\ draw) \times P(green\ on\ second\ draw)=\frac{8}{36} \times P(green\ on\ second\ draw)\]